OpenStudy (anonymous):

Express answers in simplest exact form. What is the radius of a circle with a sector area of 7 pi sq. ft. and an arc whose measure is 70°?

OpenStudy (anonymous):

Sector area means that it is a fraction of the circle. The area of a circle is \[\pi * r^2\] You can take a fraction of the circle by multiplying the area by the sector fraction. \[{70 \over 360} * \pi *r^2 = 7\] Solve for r: \[\sqrt{{{360 * 7}\over{70*\pi}}} = r\]

hero (hero):

r = 18

OpenStudy (anonymous):

thanks

hero (hero):

:)

hero (hero):

supahfart, your set up doesn't appear to be accurate

OpenStudy (anonymous):

lolz supahfart hahahaha

OpenStudy (anonymous):

You're right, I used 7, assuming that pi sq ft was a measurement XD: \[{70\over360}∗π∗r^2=7\pi\] \[\sqrt{{{360*7}\over{70}}} = r\] r = 6

OpenStudy (anonymous):

you used circumference hero. You are wrong

hero (hero):

Yes, you're right. I sped read the question. I am wrong.

OpenStudy (anonymous):

Thank you for catching my mistake though. It is greatly aprreciated

hero (hero):

yeah r =6, confirmed

hero (hero):

\[\frac {70}{360} = \frac {7\pi}{{\pi}r^2}\]